OpenStudy (ttp):

After the softball game, Alise's coach bought each of the 14 players a double-scoop or triple-scoop ice cream cone at the local ice cream shop. The coach bought a total of 37 scoops of ice cream. Use a model to determine how many girls got a triple-scoop ice cream cone. 6 7 8 9

OpenStudy (ttp):

@ngott3

OpenStudy (anonymous):

scoops total=37 #players=14 n=# of ppl for double scoops m=# of ppl for triple scoops n(2)+m(3)=37 14=(n+m) n=14-m (14-m)(2)+m(3)=37 28-2m+3m=37 m=9 therefore 9 players got the triple scoop